OpenStudy (anonymous):

Sam, our Chief Technology Officer, is in charge of ensuring that Brilliant has enough programmers to ensure help code up all these nifty features (like that "Solved by" box to the side). Bradan, our Human Resource Manager, is in charge of collecting resumes from prospective job applicants. From experience, he knows that out of every 54 resumes, he finds 6 candidates to interview. Out of every 16 interviews, he finds 2 good programmers to hire. How many resumes does Bradan need to collect, in order for Sam to find 3 good programmers to hire?

OpenStudy (anonymous):

216

OpenStudy (anonymous):

Solution : 1/9 of the curriculum collected are interviewed, and only 1/8 of these are good progammers. So, (1/9).(1/8)=1/72 of the applicants are good programmers. To find 3 of these, Bradan will have to collect 72*3=216 curriculum in order to get 3 good programmers.

Directrix (directrix):

@killua_vongoladecimo Code of conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.